User name R. Michaels/Alex
Log entry time 13:36:11 on April 1,2007
Entry number 197712
This entry is a followup to: 197698
keyword=RT & HV fixes
Alex fixed the problems in halog 107698.
- Rebooted HV crate and IOC.
- Put delay for S1 TDC data AFTER the point where it goes to trigger.